OpenStudy (anonymous):

Among all the rectangles whose perimeters are 100 feet, find the dimensions of the one with the maximum area. Please explain step by step!

OpenStudy (anonymous):

Let x be the length and y be the width of the rectangle. Perimeter = 2x + 2y 100 = 2x + 2y 50 = x + y y = 50 - x Area = xy A = x(50 - x) A = 50x - x^2 Take the derivative; set it equal to zero: A ' = 50 - 2x 0 = 50 - 2x 2x = 50 x = 25 y = 50 - x y = 50 - 25 y = 25 Thus the dimensions are (25, 25)

OpenStudy (anonymous):

Okay, thanks!